OpenStudy (anonymous):

Someone please explain the margin error formula better. My class just gave it to me and said to practice. 1.96X Sq root((P(1-p))/N)

OpenStudy (amistre64):

would you agree that spread about the mean provides us with a range of options to choose from?

OpenStudy (amistre64):



OpenStudy (amistre64):

if we wanted to find the spread about the mean that is some k% ... we would want to determine the z score related to that k% and find the values x0 and x1 that give us the boundaries of the spread.

OpenStudy (amistre64):

therefore, using the zscore formula we get:\[\large z_{k\%}=\frac{x-p}{\sqrt{pq/n}}\] \[\large z_{k\%}\sqrt{\left(\frac{pq}{n}\right)}=|x-p|\]

OpenStudy (anonymous):

But were do I put what?

OpenStudy (amistre64):

well, they give you z_k% as 1.96 (95% about the mean if memory serves) the p,q,n parts deal with the specific proportion that you are playing with.

OpenStudy (amistre64):

spose you take a survey and 6 out of 10 people like cheese. the proportion, p, is therefore 6/10. q is the complement of p, q=(1-p) = 4/10 and of course the number of people surveyed, n, is 10